OpenStudy (anonymous):

A number added to itself seven additional times is fourteen more than the number. What is the number?

OpenStudy (anonymous):

7x = x + 14

OpenStudy (anonymous):

I got 7/3

OpenStudy (anonymous):

Then that's your answer.

OpenStudy (anonymous):

but that's not in the answer choice the answer choices are a.9/14 b.2 c.7 d.-2

OpenStudy (anonymous):

hmm x * 7 - 14 = x

OpenStudy (anonymous):

Sorry, my equation was off.

OpenStudy (anonymous):

it gives me the same answer

OpenStudy (anonymous):

hrmmm

OpenStudy (anonymous):

A number added to itself seven additional times is fourteen more than the number. What is the number? If x multiplied by seven is fourteen more than x, then 7x = x + 14 hmmmmmm

OpenStudy (anonymous):

@Muzzack

OpenStudy (muzzack):

very easy x=7

OpenStudy (anonymous):

How? o.o

OpenStudy (anonymous):

7 isn't right neither my teacher said it was wrong

OpenStudy (anonymous):

Using POE, it must be 9/14.

OpenStudy (anonymous):

what is POE

OpenStudy (anonymous):

Process of elimination, but you know what? All of those answers are wrong.

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

A number added to itself 7 additional times? so if it is 2 then it would be multiplied by 9?

OpenStudy (anonymous):

Nope, still doesn't matter. This question is messed up.

OpenStudy (anonymous):

I don't know how but the answer was -2

OpenStudy (anonymous):

Not possible any way I put it. confuzzled to the max. @hartnn explain! :D

OpenStudy (anonymous):

n^7=n-14 -14/7=-2 :D